What is the percent value of R65 000 at 11% compounded annually for 4 years?

1
Expert's answer
2021-11-22T20:19:16-0500

compounded annually, n =1


"PV = \frac{FV}{(1+i)^n}"


"=\frac{R65,000}{(1+0.11)^{1×4}}"


"=\frac{65000}{(1.11)^4} =R42817.513"